OpenStudy (anonymous):

Help!! The time t (in seconds) it takes an object dropped from a height h (in feet) to reach the ground is given by the function t = 1/4 √h. HELP ME: A rock is dropped from a bridge 225 feet above the water. How long does it take the rock to reach the water?

OpenStudy (anonymous):

replace h by 225 and use a calculator

OpenStudy (anonymous):

3.75 seconds?

OpenStudy (anonymous):

\[\frac{1}{4}\times \sqrt{225}=\frac{1}{4}\times 15=3.75\] yes!

OpenStudy (anonymous):

Thank you!